Find the accumulated value of an ordinary simple annuity of Rs 3 000 per year for 7 years, if money is worth j1=8%


1
Expert's answer
2021-07-30T08:54:39-0400

P = Periodic payments

r = rate of interest

n = no of years

Future Value of Annuity "= \frac{P ( (1 + r)^n - 1 ) }{ r}"

"= \frac{3000 \times ((1 + 8\; \%)^7 - 1) }{ 8\; \%} \\

 = 26768.41"